OpenStudy (anonymous):

Sailors need to consider the speed of the wind when adjusting the sails on their boat. The force F (in pounds per square foot) on a sail when the wind is blowing perpendicular to the sail can be modeled by the function F = 0.004v2 where v is the wind speed (in knots). Find the wind speed that will produce a force of 2.5 pounds per square foot on a sail.

OpenStudy (anonymous):

somebody? @terenzreignz ?

OpenStudy (anonymous):

Any good at word problems?

terenzreignz (terenzreignz):

Well, let's have a real good think...

OpenStudy (anonymous):

I may as well be reading greek myself

OpenStudy (anonymous):

speaking of which thanks for the help on the exponents! i finally grasp them!

OpenStudy (anonymous):

ack my interenet is coming in and out! if i disapear then that's why

terenzreignz (terenzreignz):

Glad to hear it. I trust you'll excuse me for not being able to drill you on word problems, though, for two reasons... ONE, All word problems are different, there is no single unbreakable method to do them, you have to give them real thought... TWO, making up word problems takes time, not like those things I make you simplify... Having said that, let's work on this one...

OpenStudy (anonymous):

Allrgihty :)

terenzreignz (terenzreignz):

If you're given models (fancy word for equations), use them. They're probably key. We have... \[\Large F=0.004v^2\]

terenzreignz (terenzreignz):

Where F is the force, and v is the speed, aye?

OpenStudy (anonymous):

OK So that means \(\Large\ 2.5=0.004v^{2}\)

terenzreignz (terenzreignz):

Yes, exactly, and just solve for v.

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

How is your smart score NOT a hundred?

terenzreignz (terenzreignz):

Well... http://openstudy.com/users/satellite73 There's no way anybody's catching up to those stats anytime soon (btw, there can only be one user with a 100 ss)

OpenStudy (anonymous):

so v=25 or -25 right?

terenzreignz (terenzreignz):

No...

OpenStudy (anonymous):

opppsss let me try that again

OpenStudy (anonymous):

Help me here

OpenStudy (anonymous):

That sounded pushy sorry

terenzreignz (terenzreignz):

\[\Large\ 2.5=0.004v^{2}\] What to do? Well, divide both sides by 0.004.

OpenStudy (anonymous):

Ack should have thought of that

OpenStudy (anonymous):

OK so \(\LARGE\ 625=v^{2}\)

OpenStudy (anonymous):

Then what do I do?

terenzreignz (terenzreignz):

You have \[\Large v^2 = 625\] So, it's the square of v, so to get v, what do you do?

OpenStudy (anonymous):

So that would be 25 right?

terenzreignz (terenzreignz):

Yes.

OpenStudy (anonymous):

\(\LARGE\ 25^{2}\) is 625

OpenStudy (anonymous):

On these questions do I have to have two answers? Becusae I know on some questions similer to this it's like x= # or #

terenzreignz (terenzreignz):

Well, what else other than 25 would yield 625 when squared?

OpenStudy (anonymous):

well -25^2= -625

OpenStudy (anonymous):

But I don't think that would be a vaild answer?

terenzreignz (terenzreignz):

True that \[\Large -25^2 = -625\] But... \[\large (-25)^2= (-25)(-25) = (-1)(25)(-1)(25)=(-1)(-1)(25)(25)\\\large =(25)(25)= 625\]

OpenStudy (anonymous):

Wait...You lost me!

terenzreignz (terenzreignz):

/sigh/ WHAT is a negative times a negative?

OpenStudy (anonymous):

OHHH! Sorry positive :P

OpenStudy (anonymous):

all those parenthesis and multiplication caught me off guard sorry

terenzreignz (terenzreignz):

Riiight. And WHAT is \[\Large (-25)^2 = (-25)(-25) = \color{red}?\]

OpenStudy (anonymous):

625!

terenzreignz (terenzreignz):

We can't have this... when you're in elementary algebra, I (and I'm pretty sure your instructor, too) \(\large \color{red}{expect}\) you to have firm grasp of your pre-algebra... particularly dealing with negatives.

terenzreignz (terenzreignz):

And yes, the 'other' answer is -25 SINCE -25 times -25 is POSITIVE 625.

OpenStudy (anonymous):

My mistake! I promise not to make it more then once errr atleast twice :)

terenzreignz (terenzreignz):

You mean... you promise you won't do it AGAIN. EVER. okay? ^_^

OpenStudy (anonymous):

Thanks TJ :) Slow learner over here :P

terenzreignz (terenzreignz):

Can I trust you to not make that mistake again? ^_^

OpenStudy (anonymous):

Ha all I can do is try :)

OpenStudy (anonymous):

and practice!

terenzreignz (terenzreignz):

LOL The correct answer is no. You WILL make that mistake again... (because you're human) So the best we can do is try to prevent them. That's why you will do these for me... Now. Solve for the unknown: \[\Large 72 = 2\color{red}x^2\] \[\Large 78 = 3\color{blue}y^2+3\]

OpenStudy (anonymous):

Allrighty :) One second

terenzreignz (terenzreignz):

...and how old are you again?

OpenStudy (anonymous):

algebra 1

terenzreignz (terenzreignz):

That's not an age I know of... -.-

OpenStudy (anonymous):

OK the first equation would be x=6 or -6

terenzreignz (terenzreignz):

That's good. Now the second.

OpenStudy (anonymous):

working on it one second

OpenStudy (anonymous):

y=5 or -5

terenzreignz (terenzreignz):

Okay... I accept your apology. ^_^

OpenStudy (anonymous):

Great :D

terenzreignz (terenzreignz):

Let me just make sure you know what you're doing... Try this one: \[\Large -7v^2 =343\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

x=7 or -7

terenzreignz (terenzreignz):

Nope.

terenzreignz (terenzreignz):

Let's try v=7, shall we? \[\Large -7(\color{red}7)^2\]\[\Large = -7(49)\]\[\Large = -343\]\[\Large \color{red}\ne 343\]

OpenStudy (anonymous):

Oh OK the negative threw me off let me try once more

terenzreignz (terenzreignz):

Just remember to be more careful... there is no 'try once more' when exams come.

OpenStudy (anonymous):

That's the truth OK I need help

terenzreignz (terenzreignz):

Okay. Pay attention:

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

\[\Large -7v^2 =343\]

terenzreignz (terenzreignz):

\[\Large \color{blue}{-7}v^2 =343\] THIS is negative.

terenzreignz (terenzreignz):

\[\Large -7\color{red}{v^2} =343\]THIS is positive.

terenzreignz (terenzreignz):

If we multiply them: \[\Large \color{blue}{-7}\color{red}{v^2} =343\] Their product should be negative! \[\Large \color{blue}{-7v^2} =343\]

terenzreignz (terenzreignz):

Everything understood?

OpenStudy (anonymous):

Yup :) I think

terenzreignz (terenzreignz):

Remember that a square is always positive, okay (unless it's the square of zero)

OpenStudy (anonymous):

so -7*v^2 =-343

terenzreignz (terenzreignz):

Okay?

terenzreignz (terenzreignz):

NO! Listen first... and just answer when I ask you.

OpenStudy (anonymous):

Ok

terenzreignz (terenzreignz):

Okay, so you understand why \[\Large -7\color{red}{v^2} =343\]this part is always positive, right?

OpenStudy (anonymous):

right

terenzreignz (terenzreignz):

So, since -7 is negative and v^2 is positive, their product -7v^2 should be negative, right?

OpenStudy (anonymous):

Right

terenzreignz (terenzreignz):

So we know that this part \[\Large \color{blue}{-7v^2} =343\] is negative, so ask yourself...

OpenStudy (anonymous):

how did we get a positive?

terenzreignz (terenzreignz):

exactly. Why is something we're sure is negative... equal to a positive?

terenzreignz (terenzreignz):

We know that -7v^2 is negative, and yet it's made to be equal to 343, a positive.

terenzreignz (terenzreignz):

Then there are NO SOLUTIONS.

OpenStudy (anonymous):

TRICK QUESTION

OpenStudy (anonymous):

Wait right?

terenzreignz (terenzreignz):

Any questions?

terenzreignz (terenzreignz):

Okay, quick drill... what is \[\Large -8^2 = \color{red}?\]

OpenStudy (anonymous):

no...sorry there was no answer to that question! I keep forgetting a question can have no soultions http://barfblog.com/wp-content/uploads/2013/06/homer-doh.jpg

OpenStudy (anonymous):

That would be-64

terenzreignz (terenzreignz):

Good. What about \(\Large (-4)^2 = \color{red}? \)

OpenStudy (anonymous):

-16

terenzreignz (terenzreignz):

And, as predicted, you made that mistake... again...

OpenStudy (anonymous):

what mistake?